OpenStudy (anonymous):

[WILL FAN AND MEDAL] Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

We will solve this using system of equations. -9x+4y=8 -3x-y=4 y= -3x-4 (isolate the y) -9x+4(-3x-4)=8 (plug the equation for y back into the first equation) -9x-12x-16=8 -21x-16=8 -21x=24 this leaves us with x= -8/7 Now plug x into one of the original equations and solve for y. -3(-8/7)-y=4 24/7-y=4 24/7-4=y And so y= -4/7